can someone please help me with this problem i can't figure it out.
Find the future value of the ordinary annuity. Interest is compounded annually, unless otherwise indicated.
R = $900, i=8% interest compounded semiannually for 8 years.
: can someone please help me with this problem i can't figure it out.
Find the future value of the ordinary annuity. Interest is compounded annually, unless otherwise indicated.
R = $900, i=8% interest compounded semiannually for 8 years.

Answer by checkley77(3848)

that's not an answer i can choose from the answers that i can choose from are: $18,021.23; $19,280.50; $19,642.08 or $42,142.08 can you please rework it for me or give me the formula so i can try. thank's.
-------------------------------------------------------------------------
P(1+R/T)^T*Y WHERE: P=PRINCIPLE (900), R=INTEREST RATE (.08), T=TIMES THE INTEREST IS CALCULATED EACH YEAR (2) & Y=NO. OF YEARS (8).
900(1+.08/2)^2*8
900(1.04)^16
900*1.873=$1,685.70 ans.
------------------------------------------
SORRY I CANNOT GET ANY NUMBERS CLOSE TO YOUR CHOICES EVEN STARTING WITH $9,000.
9,000(1+.08/2)^2*8
9,000(1.04)^16
9,000*1.873=$16,856.83